You have just been hired as a new management trainee by Earrings Unlimited, a distributor of earrings to various retail outlets located in shopping malls across the country. In the past, the company has done very little in the way of budgeting and at certain times of the year has experienced a shortage of cash.

Since you are well trained in budgeting, you have decided to prepare comprehensive budgets for the upcoming second quarter in order to show management the benefits that can be gained from an integrated budgeting program. To this end, you have worked with accounting and other areas to gather the information assembled below.

The company sells many styles of earrings, but all are sold for the same price—$10 per pair. Actual sales of earrings for the last three months and budgeted sales for the next six months follow

January (Actual) 200000
February (Actual) 260000
March (Actual) 400000
April (Budget) 65000
May (Budget) 100000
June (Budget) 50000
July (budget) 30000
August (Budget) 28000
September (Budget) 25000

The concentration of sales before and during May is due to Mother’s Day. Sufficient inventory should be on hand at the end of each month to supply 40% of the earrings sold in the following month.

Suppliers are paid $4 for each bracelet. One-half of a month’s purchases is paid for in the month of purchase; the other half is paid for in the following month. All sales are on credit, with no discount.. The company has found, however, that only 20% of a month’s sales are collected in the month of sale. An additional 70% is collected in the following month, and the remaining 10% is collected in the second month following sale. Bad debts have been negligible.

Monthly operating expenses for the company are given below:

Variable Expenses
Sales Commissions 4% of Sales
Fixed expenses
Advertising 200000
Rent 18000
Salaries 106000
Utilities 7000
Insurance 3000
Depreciation 14000

Insurance is paid on an annual basis, in November of each year.

The company plans to purchase $16,000 in new equipment during May and $40,000 in new equipment during June; both purchases will be for cash. The company declares dividends of $15,000 each quarter, payable in the first month of the following quarter.

Other relevant data is given below:

Cash Balance as of march 31 74000

Inventory balance as of March 31 104000

Merchandise purchases for March 200000

The company maintains a minimum cash balance of $50,000. All borrowing is done at the beginning of a month; any repayments are made at the end of a month.

The company has an agreement with a bank that allows the company to borrow the exact number amount needed at the beginning of each month. The interest rate on these loans is 1% per month and for simplicity we will assume that interest is not compounded. At the end of the quarter, the company would pay the bank all of the accured interest on the loan and as much of the loan as possible while still retaining at least $50,000 in cash.

Required:

1. Prepare a master budget for the three-month period ending June 30. Include the following detailed budgets:

a. A sales budget, by month and in total.

b. A schedule of expected cash collections from sales, by month and in total.

c. A merchandise purchases budget in units and in dollars. Show the budget by month and in total.

d. A schedule of expected cash disbursements for merchandise purchases, by month and in total.

2. A cash budget. Show the budget by month and in total. Determine any borrowing that would be needed to maintain the minimum cash balance of $50000
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Earrings Unlimited Note
Sales Budget February March April May June April- Jun Qtr.
Sales (units)       260,000.00       400,000.00         65,000.00       100,000.00         50,000.00       215,000.00 A
Sell Price                 10.00                 10.00                 10.00                 10.00                 10.00 B
Sales Budget 2,600,000.00 4,000,000.00       650,000.00 1,000,000.00       500,000.00 2,150,000.00 C=A*B
Collection Budget February March April May June April- Jun Qtr.
Sales Budget    2,600,000.00    4,000,000.00       650,000.00    1,000,000.00       500,000.00 2,150,000.00 See C
20% collected in same month       130,000.00       200,000.00       100,000.00       430,000.00 D= C*20%
70% collected in next month    2,800,000.00       455,000.00       700,000.00 3,955,000.00 E= 70% of C of previous month
10% collected in second month       260,000.00       400,000.00         65,000.00       725,000.00 F= 10% of C of previous to previous month
Expected cash collections 3,190,000.00 1,055,000.00       865,000.00 5,110,000.00 G=D+E+F
Purchase Budget April May June April- Jun Qtr. July
Sales (units)         65,000.00       100,000.00         50,000.00       215,000.00 30,000.00 See A
Add: Desired Ending Inventory         40,000.00         20,000.00         12,000.00 H= 40% of A of next month
Total bracelet needed       105,000.00       120,000.00         62,000.00 I= A+H
Less: Expected Beginning Inventory         26,000.00         40,000.00         20,000.00 J= 40% of A
Bracelet Purchase Budget         79,000.00         80,000.00         42,000.00       201,000.00 K=I-J
Payment Budget March April May June April- Jun Qtr.
Bracelet Purchase Budget         79,000.00         80,000.00         42,000.00       201,000.00 See K
Purchase price                   4.00                   4.00                   4.00 L
Purchase cost       200,000.00       316,000.00       320,000.00       168,000.00       804,000.00 M=K*L
50% paid in same month       158,000.00       160,000.00         84,000.00       402,000.00 N= M*50%
50% paid in next month       100,000.00       158,000.00       160,000.00       418,000.00 O= 50% of M of previous month
Cash disbursements for merchandise purchases       258,000.00       318,000.00       244,000.00       820,000.00 P=N+O
Sales Commission Budget April May June April- Jun Qtr.
Sales Budget       650,000.00    1,000,000.00       500,000.00 2,150,000.00 See C
Sales Commission @ 4%         26,000.00         40,000.00         20,000.00         86,000.00 Q= C*4%
Cash budget April May June April- Jun Qtr.
Expected cash collections    3,190,000.00    1,055,000.00       865,000.00 5,110,000.00 See P
Total Cash Receipts 3,190,000.00 1,055,000.00       865,000.00
Estimated cash payments for
Merchandise purchases       258,000.00       318,000.00       244,000.00       820,000.00
Sales Commission         26,000.00         40,000.00         20,000.00         86,000.00
Advertising       200,000.00       200,000.00       200,000.00       600,000.00
Rent         18,000.00         18,000.00         18,000.00         54,000.00
Wages and Salaries       106,000.00       106,000.00       106,000.00       318,000.00
Utilities            7,000.00            7,000.00            7,000.00         21,000.00
New equipment         16,000.00         40,000.00         56,000.00
Dividends         15,000.00         15,000.00
Total cash payments       630,000.00       705,000.00       635,000.00 1,970,000.00
Cash increase or (decrease)    2,560,000.00       350,000.00       230,000.00
Cash Balance at the beginning of month         74,000.00    2,634,000.00    2,984,000.00
Cash Balance at the end of month 2,634,000.00 2,984,000.00 3,214,000.00
As cash Balance at the end of month is more than $ 50,000 so borrowing is not required.
Depreciation is a non cash expense so not considered.
Insurance is paid on an annual basis, in November of each year so not considered.